OpenStudy (anonymous):

a pump can throw a stream of water up to 29.6m. What is the initial speed of the stream when it leaves a hose nozzle?

OpenStudy (anonymous):

is it in upward direction??

OpenStudy (anonymous):

It doesn't specify, this is that is where I am stuck :/

OpenStudy (anonymous):

i think it is because there is no other way to find initial velocity..

OpenStudy (anonymous):

initial velocity=Vi=? final velocity=Vf=0m/s height=h= 29.6m and for upward motion g= -9.8m/s^2 using third equation of motion 2gh=Vf^2 - Vi^2 2(-9.8)(29.6)= (0)^2 - (Vi)^2 -580.16= -Vi^2 Vi= sqrt (580.16) Vi= 24.09 m/s